OpenStudy (anonymous):

^DEF is a right triangle. If FE = 8 and DE = 6, find DF.

OpenStudy (anonymous):



OpenStudy (anonymous):

@Yttrium @Jack1

OpenStudy (yttrium):

Pythagorean theorem. c^2 = a^2 + b^2 :))

OpenStudy (anonymous):

so 8 and 6 can be either a or b . im just trying to find the hypotenuse right???

OpenStudy (yttrium):

yes! :)

OpenStudy (anonymous):

ok so 8^2+6^2=c^2 64+36=c^2 100 do i sqrt it ??

OpenStudy (yttrium):

of course :))

OpenStudy (anonymous):

10??

OpenStudy (yttrium):

Yes!

OpenStudy (anonymous):

thnkkss